OpenStudy (anonymous):

Can somebody help me solve z^2/z+4-16/z+4. I really want to know how to do it so i would appreciate if someone could walk me through it! im desperate! ill give medals

OpenStudy (anonymous):

\[\frac{ Z ^{2} }{ z+4 }-\frac{ 16 }{ z+4 } \] is the equation in a form easier to read

OpenStudy (anonymous):

@Brostep0s hey could you help me with this one too?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

\[\Large\rm \frac{z^2}{z+4}-\frac{16}{z+4}\]Hmmm, our fractions have the same base, so we can combine them, yes?\[\Large\rm =\frac{z^2-16}{z+4}\]

zepdrix (zepdrix):

Here is something important that we'll need to utilize. `The difference of squares is the product of their conjugates`. It looks like this:\[\Large\rm a^2-b^2=(a-b)(a+b)\]Look familiar? :)

zepdrix (zepdrix):

In our problem, we don't have two squares. We have a z^2 and a 16. We would like to rewrite the 16 as "something squared" so we can apply this rule.

OpenStudy (anonymous):

I totally understand the whole problem now. I feel stupid lol I was looking for something I didnt need

zepdrix (zepdrix):

Oh good c: